Your project has an EV of $24,000, PV of $29,000, and AC of $45,000. If BAC is $96,000, what percent of the work has been completed?

a. 30 percent

b. 25 percent

c. 40 percent

d. 20 percent
Answer #1

b. 25 Percent

Percent Complete = (EV/BAC)*100

= (24000/96000)*100

= 0.25*100

= 25%